OpenStudy (anonymous):

7-5 exponential logarithmic equatiions. solve 5^2x=20

OpenStudy (anonymous):

.... help

OpenStudy (valpey):

\[5^{2x}=20\]Take log of both sides\[\log{5^{2x}}=\log{20}\]\[2x\log{5}=\log{20}\]\[x=\frac{\log{20}}{2\log{5}}\]You can reduce further if you want to factor out down to just the logs of prime numbers.

OpenStudy (anonymous):

But the doesn't the 2 stay with the X and at the and you dive it by 2. if you reduce it, it would be 2x=1.8613..... 1.8613.../2 x=0.930..

OpenStudy (anonymous):

it is the same thing

OpenStudy (anonymous):

okay thank you so much